Circular No. 7792

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1550–564 J. Swank, E. Smith, and C. Markwardt, Goddard Space Flight Center,

report: “Rossi X-Ray Timing Explorer observations indicate that XTE J1550–564 is again at a flux of 50–100 mCrab, which it reached in 2001 Jan. for ∼ 3 weeks. Since 2001 Oct., the ASM has indicated intermittent levels of a few ASM counts/s. In a pointed observation on 2002 Jan. 10.875 UT, the PCA confirmed a level 60 mCrab, with a very hard spectrum (power-law index 1.4 ± 0.2, cutoff 30–60 keV), corresponding to a 2–50keV luminosity of 3 × 10^{36} erg/s at 2500 pc. The spectrum includes Fe K-shell line emission, possibly in a reflection component, and ∼ 10 percent of the luminosity is in a component attributable to a disk. High-amplitude fluctuations (33-percent rms) are again present, with characteristic times of 20–100 s (cf. IAUC 7575). Based on recent behavior, a full outburst is not likely; nevertheless, additional RXTE observations are scheduled and coordinated observations at other wavelengths are encouraged.”

C. Bailyn, Yale University, adds: “YALO observations on Jan. 12.319 and 13.337 UT yield V = 18.36 and 18.33, respectively. These values are near the peak optical flux from the last two outbursts of this source.”

SUPERNOVAE 2001id, 2001ih, AND 2001ii R. Chornock, W. D. Li, and A. V. Filippenko, University of Califor-

nia, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Jan. 14 UT with the Shane 3-m reflector at Lick Observatory, reveal that SN 2001id (IAUC 7770) is a normal type-II supernova; hydrogen Balmer, Na I D, and Fe II lines are present with P-Cyg profiles. SN 2001ii (IAUC 7776) is a supernova of type Ic. SN 2001ih (IAUC 7776) is probably a type-IIb supernova; although the spectrum is somewhat noisy, it is similar to that of SN 1996cb, ∼ 1-2 months past maximum light (Qiu et al. 1999, A.J. 117, 736). Well-developed He I lines are present, as well as an absorption line identified with Hα.

COMET P/2001 Y1 (HELIN-LAWRENCE) Further to IAUC 7790, K. Sárneczky reports that his 300-s unfiltered

CCD images taken on Jan. 11.2 UT show a diffuse, 8′′ coma and a faint, narrow, 13′′ tail in p.a. 283^{o}.

2002 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green